# Skip to the Good Part: Representation Structure & Inference‑Time Layer Skipping in Diffusion vs. Autoregressive LLMs

### 08 Mar 2026 v1

@@ 3 Layer-skip at Inference
11:  if $l\notin\mathcal{L}_{\text{skip}}$ and $l-1\notin\mathcal{L}_{\text{skip}}$ and $l+1\notin\mathcal{L}_{\text{skip}}$ and $\mathcal{S}[l]\geq\tau$ then
@@
13:   if $|\mathcal{L}_{\text{skip}}|\geq n_{\text{max}}$ then
@@
4 Experimental Setup
Models.
We evaluate three families to disentangle training objective and initialization: (i) a native diffusion LLM, LLaDA (we use the 8B Base / Instruct checkpoints) (Nie et al., 2025); (ii) a native autoregressive (AR) model, Qwen2.5 (7B Base / Instruct) (Yang et al., 2024); and (iii) an AR-initialized diffusion LLM, Dream-7B (Instruct) (Ye et al., 2025). Unless stated otherwise, all models are evaluated with their public inference code and default tokenizers.
Benchmarks.
Prompting and answer extraction.
For GSM8K, we use a few-shot rationale prompt with an explicit Final Answer: line; we strip formatting and compare normalized numbers. For HumanEval and MBPP, we request a single Python function and evaluate with the official test suites; we report pass@1 and pass@k. The same prompts are used across models to ensure comparability.We follow exact inference setting of (Chen & Liu, 2026)
Decoding & sampling.
[…]
Layer-skipping evaluation (ours).
5 Results
Native dLLMs Enable Aggressive Layer Skipping. Table 1 and Figure 1 show that native diffusion LLMs are markedly more robust to layer skipping than autoregressive models. For LLaDA, skipping 6 layers (18.75% FLOPs reduction) preserves 88.2–102.1% of baseline performance across all evaluated tasks. Even at an 8-layer skip (25% FLOPs reduction), performance retention remains high (62.7–91.8%), placing LLaDA firmly in the favorable efficiency–quality regime (top right of Figure 1).
In contrast, autoregressive models degrade rapidly under similar interventions. Qwen2.5 experiences severe performance collapse: skipping only 2 layers (7.14% FLOPs reduction) reduces retention to 34.9–75.3%, indicating that AR representations are tightly coupled and lack depth-wise redundancy. Notably, Dream‑7B—despite diffusion fine‑tuning—exhibits similar brittleness. At a 2-layer skip, Dream‑7B achieves only 60.5–81.4% retention, closely tracking Qwen2.5 and substantially underperforming LLaDA (100–123.1%). This provides strong evidence that autoregressive pre-training induces a persistent representational structure that diffusion objectives alone do not override, preventing the emergence of the coarse‑to‑fine hierarchy observed in native dLLMs.
[…]
@@ 5 Results
Table 1: Performance comparison across models and layer skipping configurations. Values represent retention percentages relative to baseline (0 layers skipped), with absolute accuracy shown in parentheses for baseline rows. Missing entries (-) indicate performance below retention threshold. Higher values ($\uparrow$) indicate better retention.
Benchmark Layers Skipped Llada-Instruct Dream-Base Dream-Instruct Qwen2.5
GSM8K 0 (baseline) 100.0 (0.83) 100.0 (0.77) 100.0 (0.81) 100.0 (0.49)
2 101.3 76.8 84.6 34.9
3 96.7 62.6 71.2 15.9
4 102.5 44.5 25.0 20.6
5 97.7 - - -
6 91.8 - - -
8 91.8 - - -
MATH500 0 (baseline) 100.0 (0.37) 100.0 (0.34) 100.0 (0.43) 100.0 (0.18)
2 108.5 81.4 78.2 -
3 89.4 60.5 54.6 -
4 89.4 58.8 32.7 -
5 93.6 - - -
6 102.1 - - -
8 70.2 - - -
HumanEval 0 (baseline) 100.0 (0.51) 100.0 (0.65) 100.0 (0.64) 100.0 (0.68)
2 100.0 66.2 70.3 64.7
3 100.0 63.1 59.4 48.5
4 92.2 43.1 26.6 41.2
5 82.3 - - -
6 88.2 - - -
8 62.7 - - -
MBPP 0 (baseline) 100.0 (0.52) 100.0 (0.43) 100.0 (0.71) 100.0 (0.81)
2 123.1 60.5 70.3 75.3
3 115.4 74.4 54.1 55.6
4 115.4 53.5 37.8 35.8
5 111.5 - - -
6 94.2 - - -
8 71.1 - - -
5.1 Layer Distribution and Skip Sensitivity
Tables 2 and 3 reveal that consecutive layer skipping is catastrophic. For LLaDA at 6-layer skip, allowing consecutive removal drops GSM8K retention from 91.8% to 75.3% and HumanEval from 88.2% to 64.7%. Dream-7B shows extreme sensitivity: 2-layer consecutive skip collapses GSM8K retention from 76.8% to 14.1%.
Our algorithm avoids this by maintaining representational continuity. Analysis and Fig.8 shows skipped layers concentrate in early network depth (first 40–60%), aligning with our observation that early layers establish coarse representations with high redundancy, while later layers perform critical fine-grained refinement.
Table 2: Performance retention for LLada model with layer-skip when allowed and not allowed to skip consecutive layers. Higher ($\uparrow$) the better
Layers Skipped GSM8k HumanEval
Not Allowed Allowed Not Allowed Allowed
4 102.5 95.1 92.2 88.2
6 91.8 75.3 88.2 64.7
8 91.8 45.8 62.7 23.5
Table 3: Performance retention for Dream model with layer-skip when allowed and not allowed to skip consecutive layers. Higher ($\uparrow$) the better
Layers Skipped GSM8k HumanEval
Not Allowed Allowed Not Allowed Allowed
2 76.8 14.1 78.2 67.7
3 62.6 18.2 54.6 53.8
4 44.5 16.2 32.7 23.1
Figure 8: Which layers are skipped? Empirical distribution over layer indices selected for skipping by Algorithm 1on LLaDA, aggregated across evaluation prompts on GSM8k.The plotted value at layer $i$ is the fraction (probability) of runs/examples in which layer $i$ is chosen to be skipped. Left: consecutive (adjacent) layers may be skipped, allowing contiguous skip blocks. Right: consecutive skipping is disallowed, restricting selections to non-adjacent layers.
6 Related Work
Diffusion Language Models and Representations. Diffusion language models (dLLMs) replace autoregressive decoding with bidirectional denoising objectives, enabling parallel decoding and global context modeling. Foundational work on discrete diffusion (Austin et al., 2021) led to recent dLLMs such as SEDD (Lou et al., 2023), and LLaDA (Nie et al., 2025; Bie et al., 2025), which achieve competitive language modeling performance. Dream‑7B (Ye et al., 2025) adapts pretrained AR models to diffusion training, while MDLM (Sahoo et al., 2024) simplifies diffusion objectives. Despite this progress, how diffusion objectives shape internal representations—especially relative to AR and AR‑initialized models—remains insufficiently understood.
Additionally, efforts have been made to integrate KV caching mechanisms to reduce redundant computation (Ma et al., 2025; Liu et al., 2025). An alternate line of work focuses on step distillation of dLLMs (Deschenaux & Gulcehre, 2025; Qian et al., 2026) towards acceleraing dLLM inference.
Representational Structure and Initialization Effects. Prior work shows that language models organize representations hierarchically across depth, with earlier layers capturing coarse features and deeper layers refining task‑specific abstractions (Jawahar et al., 2019). While representational dynamics in AR models have been studied, systematic analyses comparing AR and diffusion models are rare. Our work provides direct, layer‑ and token‑level evidence of this initialization bias in AR‑adapted dLLMs and contrasts it with the representational redundancy that emerges in native diffusion models.
7 Summary and Future Work
We investigated how training objectives shape internal representations in language models by comparing native diffusion LLMs (LLaDA), autoregressive models (Qwen2.5), and AR-adapted diffusion models (Dream-7B). Our analysis reveals that: (1) AR-adapted dLLMs retain AR-like representations despite diffusion training, demonstrating strong initialization bias; (2) diffusion objectives reduce recency bias and promote global representations. (3) native dLLMs exhibit coarse-to-fine hierarchies with early-layer redundancy.
Leveraging these insights, we introduced a static layer-skipping strategy that achieves substantial FLOPs reduction with minimal performance loss on native dLLMs, while AR models prove more brittle—validating that objective-induced redundancy enables efficient inference.
Future Work: Several directions are worth exploring. First, dynamic or input‑adaptive skip policies may offer further gains by allocating computation based on difficulty. Second, combining layer skipping with complementary efficiency techniques such as KV caching could yield compounded speedups. Third, lightweight fine‑tuning with adapters may enable more aggressive skipping or pruning while preserving quality. Finally, extending our representational analysis and skipping strategies to multimodal diffusion architectures remains an exciting avenue for future research.
8 Impact Statement
This work proposes inference-time layer skipping for diffusion language models to improve computational efficiency. By reducing inference-time FLOPs, our approach can lower hardware and energy costs, helping broaden access to large language models and supporting more environmentally sustainable deployment. Our representational analyses also contribute tools for better understanding model internals, which may aid future work on interpretability and robustness.
Our finding that AR-adapted diffusion models (Dream-7B) retain strong representational signatures from their AR initialization—despite substantial diffusion training—has important implications for model adaptation practices. This persistent initialization bias suggests that safety properties, biases, or failure modes from pre-trained models may transfer to fine-tuned variants in non-obvious ways. Practitioners adapting models to new objectives should carefully validate behavioral properties rather than assuming the new training objective fully overrides initialization effects.
As with any efficiency optimization, skipped computations could affect behaviors not covered by standard benchmarks, particularly in safety-critical or out-of-distribution settings. We recommend domain-specific validation when deploying the method in high-stakes applications. Overall, we view this work as a technical contribution intended to improve the practicality and accessibility of modern language models.

### 27 Apr 2026 v2

@@ Appendix C Layer-skip Algorithm
11:  if $l\notin\mathcal{L}_{\text{skip}}$ and $l-1\notin\mathcal{L}_{\text{skip}}$ and $l+1\notin\mathcal{L}_{\text{skip}}$ then
@@
13:   if $|\mathcal{L}_{\text{skip}}|\geq n_{\text{skip}}$ then
@@
Appendix D Additional Results
We provide accuracy retention for different numbers of layers skipped for all the models considered in experiments in Table 3
@@ Appendix D Additional Results
Table 3: Model Performance Retention Rate Across Layer Skipping. Values represent retention percentages relative to baseline (0 layers skipped), with absolute accuracy shown in parentheses for baseline rows.
Task Layers Skipped Models
LLada-8B-Instruct Dream-7B-base Dream-7B-Instruct Qwen2.5-7B
GSM8K 0 (baseline) 100% (0.79) 100% (0.74) 100% (0.80) 100% (0.21)
1 101% (0.80) 99% (0.73) 99% (0.79) 110% (0.23)
2 95% (0.75) 101% (0.75) 91% (0.73) 48% (0.10)
3 100% (0.79) 77% (0.57) 71% (0.57) 5% (0.01)
4 94% (0.74) 68% (0.50) 18% (0.14) 24% (0.05)
5 94% (0.74) 55% (0.41) 4% (0.03) 19% (0.04)
6 94% (0.74) 22% (0.16) 4% (0.03) -
7 104% (0.82) 1% (0.01) 0% (0.00) -
8 94% (0.74) - - -
MATH500 0 (baseline) 100% (0.38) 100% (0.36) 100% (0.48) 100% (0.07)
1 100% (0.38) 92% (0.33) 73% (0.35) 114% (0.08)
2 108% (0.41) 69% (0.25) 77% (0.37) 143% (0.10)
3 92% (0.35) 53% (0.19) 50% (0.24) 71% (0.05)
4 87% (0.33) 50% (0.18) 31% (0.15) 100% (0.07)
5 105% (0.40) 33% (0.12) 4% (0.02) 57% (0.04)
6 97% (0.37) 25% (0.09) 0% (0.00) 0% (0.00)
7 84% (0.32) 0% (0.00) 0% (0.00) -
8 66% (0.25) - - -
HumanEval 0 (baseline) 100% (0.53) 100% (0.64) 100% (0.62) 100% (0.46)
1 100% (0.53) 84% (0.54) 76% (0.47) 98% (0.45)
2 96% (0.51) 95% (0.61) 90% (0.56) 78% (0.36)
3 100% (0.53) 66% (0.42) 73% (0.45) 87% (0.40)
4 91% (0.48) 61% (0.39) 29% (0.18) 33% (0.15)
5 77% (0.41) 39% (0.25) 5% (0.03) 11% (0.05)
6 83% (0.44) 8% (0.05) 0% (0.00) 7% (0.03)
7 79% (0.42) 0% (0.00) 0% (0.00) -
8 60% (0.32) - - -
MBPP 0 (baseline) 100% (0.56) 100% (0.44) 100% (0.75) 100% (0.63)
1 107% (0.60) 66% (0.29) 96% (0.72) 76% (0.48)
2 107% (0.60) 48% (0.21) 80% (0.60) 62% (0.39)
3 97% (0.54) 91% (0.40) 68% (0.51) 51% (0.32)
4 100% (0.56) 80% (0.35) 37% (0.28) 27% (0.17)
5 97% (0.54) 32% (0.14) 11% (0.08) 34% (0.21)
6 93% (0.52) 14% (0.06) 12% (0.09) 19% (0.12)
7 89% (0.50) 3% (0.01) 0% (0.00) -
8 77% (0.43) - - -
We show wall-time savings for LLada-8B-Instruct model is Table 4 showing that our method results in faster inference with minimal accuracy drop. Furthermore, we integrate our method with dual-cache in Table 5 and show that accuracy retention is still 88% when skipping 4 layers, that is, 12.5% FLOPs saving
Table 4: LLada-8B-Instruct Performance and Computation Savings on various downstream tasks
Task Metric Layers Skipped
0 (baseline) 2 4 6 8
GSM8K Accuracy retention 100% 95% 94% 94% 94%
Wall time reduction 0% -8% 2% 4% 10%
MATH500 Accuracy retention 100% 108% 87% 97% 66%
Wall time reduction 0% -3% 2% 10% 29%
HumanEval Accuracy retention 100% 96% 91% 83% 60%
Wall time reduction 0% 0% 17% 21% 27%
MBPP Accuracy retention 100% 107% 100% 93% 77%
Wall time reduction 0% 12% 16% 25% 29%
Table 5: Llada-8B-Instruct evaluated on GSM8K with Layer Skipping and Dual Cache Method
Layers Skipped Retention Score
0 (baseline) 100 (0.78)
2 104 (0.81)
3 101 (0.79)
4 88 (0.69)
5 77 (0.60)
6 68 (0.53)
Appendix E Additional Analysis and Visualizations
E.1 Detailed Token-wise Similarity Analysis
We provide comprehensive token-wise similarity visualizations to complement the layer-wise analysis presented earlier. These reveal how hidden state representations evolve across tokens within individual layers, providing deeper insight into the recency bias and global vs. local representation patterns discussed in the main text.
Figure 8 shows token-wise cosine similarity across all 32 layers of LLaDA. Early layers (0–15) exhibit consistently high similarity (>0.9>0.9) between consecutive tokens, indicating smooth representational transitions with minimal recency bias. This validates our hypothesis that native dLLMs establish stable global context in early layers. Later layers (16–31) show increased variability and lower similarity, reflecting task-specific refinement and decoder-like behavior where representations are actively updated for generation.
In stark contrast, from Fig.  9 reveals that Dream-7B maintains significant recency bias across all layers. Consecutive token representations show substantial changes throughout network depth, mirroring the incremental token-by-token refinement characteristic of autoregressive models. This pattern persists despite diffusion training, providing mechanistic evidence that AR initialization creates persistent representational structure. The lack of hierarchical abstraction—with similar update patterns across all depths—explains Dream-7B’s brittleness under layer skipping (Table 1), where it behaves more like Qwen2.5 than LLaDA.
E.2 Layer-wise Token Similarity by Depth
Figures below show token-wise similarity patterns grouped by network depth, revealing the transition from global to local representations:
(a) Token-wise similarity in early layers (0–7). LLaDA shows uniformly high similarity across all tokens, indicating stable global representations. Dream-7B and Qwen2.5 exhibit lower similarity with visible recency effects, demonstrating incremental AR-style processing even in early layers.
(b) Token-wise similarity in early-middle layers (8–15). LLaDA maintains high similarity, while Dream-7B and Qwen2.5 continue showing strong recency bias. The divergence between native dLLM and AR-initialized models becomes more pronounced.
Figure 6: Token-wise similarity across early and early-middle layers. (a) Layers 0–7 and (b) layers 8–15.
(a) Late-middle layers (16–23). LLaDA begins transitioning to lower similarity, indicating the onset of task-specific refinement. Dream-7B and Qwen2.5 maintain consistent recency patterns throughout depth.
(b) Late layers (24–31). LLaDA shows increased variability and lower similarity, reflecting active decoder-like refinement for generation. Dream-7B and Qwen2.5 continue incremental updates with strong recency bias, lacking the global transition observed in native dLLMs.
Figure 7: Token-wise similarity in later layers. (a) Late-middle (16–23) and (b) late (24–31).
Figure 8: Token-wise cosine similarity across all layers for LLaDA. Each subplot shows the cosine similarity between consecutive token representations ($\mathbf{h}_{\ell,i}$ and $\mathbf{h}_{\ell,i+1}$) within a specific layer $\ell$. High similarity indicates smooth representational transitions, while low similarity indicates significant representational changes between tokens. LLaDA exhibits consistently high token-wise similarity across early layers, demonstrating minimal recency bias and global representational abstraction. Later layers show increased variability, indicating task-specific refinement and decoder-like behavior. This pattern validates our hypothesis that native dLLMs develop coarse-to-fine abstraction hierarchies, with early layers establishing stable global context and later layers performing iterative refinement.
Figure 9: Token-wise cosine similarity across all layers for Dream-7B. Each subplot shows the cosine similarity between consecutive token representations within a specific layer. In stark contrast to LLaDA (Figure 8), Dream-7B exhibits significant recency bias across all layers, with substantial representational changes for each new token often. The lack of hierarchical abstraction—with similar token-by-token update patterns across all depths—confirms that Dream-7B retains AR-like incremental refinement and retain different representational abstraction compared to native dLLMs. This provides mechanistic evidence for the initialization bias observed in our layer-skip experiments (Table 1), where Dream-7B’s brittleness mirrors Qwen2.5 despite diffusion training.